OpenStudy (anonymous):

Find the Integral lnx/3x

OpenStudy (anonymous):

\[\int\limits_{}^{}(lnx/3x)dx\]

OpenStudy (anonymous):

by parts comes to mind

OpenStudy (anonymous):

the 1/3 is a constant so your pretty much left with determining the ln(x)/x portion

OpenStudy (anonymous):

So I can make it \[1/3\int\limits_{}^{}lnx/x\] Would my u be x or lnx?

OpenStudy (anonymous):

hmm, u sub 1/x is the derivative of ln(x) let u = ln(x) du = dx/x, x du = dx u*x du ------ = u du x

OpenStudy (anonymous):

So xlnx/x?

OpenStudy (anonymous):

no what is the integral of u du ?

OpenStudy (anonymous):

\[let~u=ln(x)\] \[\frac 13\int u~du\]

OpenStudy (anonymous):

\[\frac 13\int\frac 1x~ln(x)~dx\] \[u = ln(x)~:~du=dx/x\\~~~~~~~~~~~~~~~~ x~du = dx\] \[\frac 13\int\frac xx~u~du\] \[\frac 13\int u~du\]

OpenStudy (anonymous):

so u^2/2, times 1/3 and we know what u is equal to